OpenStudy (anonymous):

On Halloween, Riki approaches a door in her cat costume and a plastic pumpkin basket in her hand. The adult who is handing out treats has a dish of three kinds of candy: 48 tiny chocolate bars, 24 gummy candies, and 6 taffy chews. Riki hates taffy chews. What is the probability she will get either a gummy candy or a chocolate bar? (Assume she only gets one treat randomly selected.)

OpenStudy (photon336):

Basic probability formula. \[\frac{ number~desired~outcomes }{ total~possible~outcomes }\] first ask yourself how many total candies do we have? and does the probability that choosing at choosing a gummy candy will influence the chance of getting a chocolate bar>?

OpenStudy (anonymous):

you have to add up the total amount of candy. which would be 78 pieces of candy. Then add up the amount of gummy candy and the amount of chocolate candy, which would be 72. The probability that she would get a gummy or chocolate would be 92%

OpenStudy (photon336):

\[p(B) ~Probability~of~getting~a~gummy = \frac{ number~of~gummies }{ total~piece~of~candy }\] \[p(A) ~Probability~of~getting~a~chocolate~\bar = \frac{ number~of~chocolates }{ total~piece~of~candy }\] now do these events influence each-other? does the chance of getting a chocolate bar influence my chance of getting a gummy? Well no. because it's totally random. so we can actually add both probabilities together, remember the key word here is OR \[p(A)+p(B)\] = total probability. let me know what you get @chachi123

OpenStudy (photon336):

does this make sense?